OpenStudy (moonlitfate):

A radius of a cylinder is 1/3 times its height. The height is increasing at a rate of 2 cm per minute. Determine the change of volume to the time of the cylinder when h = 6.

OpenStudy (anonymous):

The volume of a cylinder is \[V=\frac{1}{3}\pi r^2h\] You're told that the radius of the cylinder is 1/3 of the height, so you know \[r=\frac{1}{3}h,\] which gives you \[\begin{align*}V&=\frac{1}{3}\pi \left(\frac{1}{3}h\right)^2h\\ &=\frac{1}{27}\pi h^3\end{align*}\] And, you also know that \[\frac{dh}{dt}=2\frac{\text{cm}}{\text{min}}\] Try differentiating the last volume equation with respect to h.

OpenStudy (kainui):

These questions are essentially always asking you to find a formula for the quantities described and then do implicit differentiation with respect to time. @SithsAndGiggles The volume of a cylinder is plain old V=pi*r^2*h. There's no 1/3 involved.

OpenStudy (kainui):

 Wait, I still think you've gotten something wrong here.

OpenStudy (anonymous):

Let me correct myself again... \[V=\frac{1}{9}\pi h^3\] Thanks again.

OpenStudy (moonlitfate):

So, I would have to take the derivative of that formula @SithsAndGiggles? >_<; Word problems on top of implicit differentiation.... my poor brain, hah.

OpenStudy (anonymous):

Yes, you would have to do that. You shouldn't be surprised, though. Calculus concepts are everywhere in a dynamic world.

OpenStudy (moonlitfate):

Doesn't make them easy. :p

OpenStudy (moonlitfate):

So the derivative of that: \[(\pi h^2)/3\]

OpenStudy (anonymous):

Yes. You have to find the rate of change of volume (dV/dt) when h = 6. And now that I see another mistake, I should say that I meant to say "differentiate with respect to t," not h. So, you have \[V'(t)=\frac{dV}{dt}=\frac{1}{3}\pi h^2\frac{dh}{dt}\] All you need to do now is find V'(6).

OpenStudy (moonlitfate):

Would the answer be: \[24\pi\]